### v2

@@ 3.2.1 Decentralised Identity
$<did:method:method-specifier>$
@@
    $did$: The scheme indicating that it’s a decentralized identifier.
@@
    $method$ : Specifies the DID method, which defines how the identifier is created, resolved, and managed on a specific blockchain or distributed system.
@@
    $method-specifier$ : A unique string generated according to the method’s rules, often linked to a public key or hash.
@@ 4 Proposed Architecture
Figure 4: Proposed Architecture
@@ 4.2 Dispute Resolution: How to handle cheating robots?
    When the dispute occurs because of a cheating robot: If a cheating robot $R_{1}$ tries to close the state channel unilaterally with a state that was not agreed upon by both, the robot $R_{2}$ can initiate a dispute by revealing the most recent valid state. A challenge period is a time window during which one party can contest the closure or the state submitted by the other party. During the challenge period, the disputed state can be reviewed, and the honest robot can reclaim its rightful share by presenting cryptographic evidence or other proofs of the legitimate state. Moreover, punishment mechanisms can also be introduced to punish misbehaving robots in order to encourage honest behaviour.
@@ 5 Implementation
We implemented the robot clients in Typescript. The p2p communication between the clients was facilitated by the libp2p protocol. The Quorum blockchain was used as layer 1, with the smart contracts written in Solidity.
Each robot in the system was globally identified with a DID. The P2P communication between the two robots was established with libp2p multi-address. A multiaddress or multiaddr, is a convention for encoding multiple layers of addressing information into a single “future-proof” path structure. It defines human-readable and machine-optimized encodings of common transport and overlay protocols and allows many layers of addressing to be combined and used together. A multiaddr looks like as follows:
@@
We deployed four smart contracts:-
@@
Once a robot obtains the multiaddr of its peer, it sends a digitally signed off-chain transaction to its peer. A digitally signed off-chain transaction looks like the one shown in Fig. 7. The blockchain rejects an outdated transaction submitted by a misbehaving robot in order to obtain an increased credit score or energy units. Figure 9 shows a stale transaction rejected by the smart contract on the blockchain.
@@
Figure 9: A stale transaction from a cheating robot declined by the blockchain
@@ 6 Results
We evaluated the performance of RoboComm based on the average transaction time of off-chain state channel operations (TxSign, TxVerify) and DID document generation(DID Doc Gen) time. The implementation was run on a Raspberry Pi 5 with a 2.4 GHz, A-Series CPU and 8 GB of RAM. Moreover, the blockchain network was run on a MacBook Pro with an Apple M2 chip and 8GB of RAM. The Fig. 10 shows the plot of time required for signing an off-chain transaction, verifying an off-chain transaction, and DID document generation. The statistics are averaged over 1000 independent runs.
@@
We also measured the size of the robot DID document and the off-chain transaction payload as shown in table 1 .
@@
The preliminary results indicate that RoboComm offers the necessary scalability with limited storage overhead. Given the time for off-chain transactions (signature and verification) over the state channel and the generation time for the robot’s DID document (1.873 ms., 1.28 ms., and 3.046 ms., respectively), clearly demonstrates the suitability for real-world applications.
@@ 8 Related Works
Privacy and trust management in robotics are sufficiently studied in the literature. A taxonomy of privacy constructs for privacy-sensitive robotics is provided in [18] whereas [19] provides themes that should comprise privacy-sensitive robotics research. In [20] [21] authors introduce a blockchain-based framework for handling misbehaving and byzantine robots in the system. A blockchain-based approach for collaborative decision-making with a lightweight consensus mechanism is proposed in [1]. Moreover, trust management and trustworthiness in robotic systems are studied in[22] [23][24]. There have also been some works in the literature that exploit blockchain for trust and reputation management, such as [25][26]. Although blockchain-based systems are better than centralized systems, they are not scalable enough for practical deployments due to low transaction throughput and resource-intensive consensus algorithms, which are not ideal for low-resource robots such as swarm robots, which have limited processing and storage resources available.

### v3

@@ 3.2.1 Decentralised Identity
<did:method:methodspecifier><did:method:method-specifier>
@@
    diddid: The scheme indicating that it’s a decentralized identifier.
@@
    methodmethod : Specifies the DID method, which defines how the identifier is created, resolved, and managed on a specific blockchain or distributed system.
@@
    methodspecifiermethod-specifier : A unique string generated according to the method’s rules, often linked to a public key or hash.
@@ 4 Proposed Architecture
Figure 4: Architecture of RoboComm
@@ 4.2 Dispute Resolution: How to handle cheating robots?
    When the dispute occurs because of a cheating robot: If a cheating robot $R_{1}$ tries to close the state channel unilaterally with a state that is not agreed upon by both, the robot $R_{2}$ can initiate a dispute by revealing the most recent valid state. A challenge period is a time window during which one party can contest the closure or the state submitted by the other party. During the challenge period, the disputed state can be reviewed, and the honest robot can reclaim its rightful share by presenting cryptographic evidence or other proofs of the legitimate state. Moreover, punishment mechanisms can also be introduced to punish misbehaving robots in order to encourage honest behaviour.
@@ 5 Implementation
We implement the robot clients in Typescript. The p2p communication between the clients is facilitated by the libp2p protocol. The Quorum blockchain is used as layer 1, with the smart contracts written in Solidity.
Each robot in the system is globally identified with a DID. The P2P communication between the two robots is established with libp2p multi-address. A multiaddress or multiaddr, is a convention for encoding multiple layers of addressing information into a single “future-proof” path structure. It defines human-readable and machine-optimized encodings of common transport and overlay protocols and allows many layers of addressing to be combined and used together. A multiaddr looks like as follows:
@@
We deploy four smart contracts:-
@@
Once a robot obtains the multiaddr of its peer, it sends a digitally signed off-chain transaction to its peer. A digitally signed off-chain transaction looks like the one shown in Fig. 7. The blockchain rejects an outdated transaction submitted by a misbehaving robot in order to obtain an increased credit score or energy units.
@@
6 Experiment
We measure the effectiveness of RoboComm, by applying it to robots performing a collaborative payload transport task. We rely on Agent Based Modeling(ABM) for the experiment. ABM is an effective approach for modeling the behavior of swarm robots[18][19]. Using the Mesa[20] agent-based modeling framework, a swarm of robots is deployed on a structured grid and tasked with delivering payloads to clustered goal zones modeled via a Gaussian Mixture distribution. Each robot in the swarm is assigned a delivery goal. Details about the setting of parameters for the modeling can be found in Table 1.
Two configurations are compared: a baseline swarm operating without energy sharing and a RoboComm-enabled swarm where low-energy robots can receive energy transfers from nearby idle robots. Performance is assessed on 100 independent runs of 50 simulation steps, with metrics including total deliveries, stalled robots, swarm energy dynamics, and average robot energy as shown in Fig. 9.
Parameter Setting
Parameter Value/Details
No. of Robots 10
Grid Size 20 X 20
Initial Energy [10, 50] Units
Delivery Goal 5 deliveries/robot
Energy Consumption 1 unit/step
Transfer Trigger Receiver Energy $\leq$ 2 Units
Delivery Zone Placement 5 clusters, 50 points each, $\sigma$=1.5
Simulation Steps 50
Runs 100
Table 1: Experiment parameters
@@ 7 Results
We evaluate RoboComm performance based on the average transaction time of off-chain state channel operations (TxSign, TxVerify) and DID document generation time (DID Doc Gen). The implementation is run on a Raspberry Pi 5 with a 2.4 GHz, A-Series CPU and 8 GB of RAM. Moreover, the blockchain network runs on a MacBook Pro laptop with an Apple M2 chip and 8GB of RAM. Fig. 10 shows the graph of time required for signing an off-chain transaction, verifying an off-chain transaction, and DID document generation. The statistics are averaged over 1000 independent runs.
Figure 9: Agent based modeling results of swarm performance with and without RoboComm for collaborative payload transport task
@@
We also measure the size of the robot DID document and the payload of the off-chain transaction as shown in table 2 .
@@
The preliminary results indicate that RoboComm offers the necessary scalability with limited storage overhead. Millisecond range off-chain transactions (signature and verification) over the state channel and the generation time for the robot’s DID document (1.873 ms., 1.28 ms., and 3.046 ms., respectively), clearly demonstrates the feasibility of the approach for real world deployments. Further, results from the agent based modeling shows that RoboComm reduces premature robot failures, balances swarm energy distribution, and thereby increases task completion efficiency compared to the non energy sharing approach.
@@ 9 Related Works
Privacy and trust management in robotics are sufficiently studied in the literature. A taxonomy of privacy constructs for privacy-sensitive robotics is provided in [21] whereas [22] provides themes that should comprise privacy-sensitive robotics research. In [23] [24] authors introduce a blockchain-based framework for handling misbehaving and byzantine robots in the system. A blockchain-based approach for collaborative decision-making with a lightweight consensus mechanism is proposed in [1]. Moreover, trust management and trustworthiness in robotic systems are studied in[25] [26][27]. Some works in the literature have also taken advantage of blockchain for trust and reputation management, such as [28][29]. Although blockchain-based systems are better than centralized systems, they are not scalable enough for practical deployments due to low transaction throughput and resource-intensive consensus algorithms, which are not ideal for low-resource robots such as swarm robots, which have limited processing and storage resources available.